# Obfuscated Location Disclosure for Remote ID Enabled Drones

### v1

@@ 7 Proof-of-Concept Implementation
Hardware Details. We use two different devices as a drone, i.e., a general-purpose laptop and an embedded device, to evaluate the performance of OLO-RID on high-end and medium-end drones. Our first PoC uses a general-purpose laptop equipped with one CPU running at 2.60GHz and 16GB of Random Access Memory (RAM). Such features align with the resources available onboard medium-high-end drones, e.g., the DJI Matrice 300 [35]. The second PoC uses as the drone a RPI Model 3B+, featuring a 1.4GHz CPU and 1GB of RAM [36]. The resources available on the RPI align with the ones of medium-low-end drones such as the DJI Mini 2 [37]. Also, note that such a choice aligns with the methodology used by other scientific contributions in the drone security domain [19]. Note that both the laptop and the RPI integrate an IEEE 802.11b/g/n wireless module, useful for transmitting and receiving RID-compliant WiFi messages over the air.
Software. We implemented OLO-RID using the C programming language to reduce runtime delays deriving from language interpretation as much as possible. We used the well-known library MIRACL CORE to implement cryptography operations required by the ECIES encryption scheme [38]. Note that we chose MIRACL Core due to its well-known excellent performance on low-powered IoT systems and the available support for several elliptic curves, as required for running ECIES. Such a choice also aligns with relevant literature on security solutions for constrained drones [6]. We report in Tab. II the Elliptic Curves we used within the ECIES scheme. To show the difference between curves of the same security levels, we have also considered two additional curves, namely BN254 and BLS48556. Using such curves could give us information about the difference in performance between two curves with the same level of security. Note that the curve BN254 was designed for 128-bit security [39] but, recently, its actual security decreased to 100-bit [40].
@@
BN254 64 100 [40]
secp256r1 (NIST256) 64 128 [41]
secp384r1 (NIST384) 96 192 [41]
secp521r1 (NIST521) 132 256 [41]
BLS48556 140 256 [42]
@@
We ran OLO-RID on the RPI as a native process. Instead, on the general-purpose laptop, to isolate our protocol from concurrent processes as much as possible, we ran OLO-RID within a Linux-based virtual environment featuring 4 CPU cores and 8GB of RAM. The source code of our solution is available open-source at [10].
@@ 8.1 Experimental Methodology
  • RQ1: What is the runtime overhead of OLO-RID on regular drones regarding computation time, RAM usage, energy expenditure, and communication bandwidth with various security levels?
@@
  • RQ2: How can we improve the location privacy provided to the drone at runtime, and how do such improvements affect the overhead of our solution?
@@
  • RQ3: What is the utility of the location disclosed through OLO-RID by the drone across various configuration parameters while considering our reference use cases discussed in Sec. 3.3?
@@
UAVs are constrained devices, characterized by (one or more) constraints regarding available computational, memory, communication bandwidth, and energy resources. Such constraints make it challenging to implement and run a suitable technique for location obfuscation while still creating and broadcasting RID messages within the time limitations imposed by the RID regulation ($1$ second). RQ1 sets a baseline for the overhead of our solution when using different hardware and various ECs. RQ2 investigates the impact in terms of location privacy and overhead on the drone derived by changing relevant location obfuscation parameters (i.e., $\epsilon$). Finally, RQ3 evaluates OLO-RID in the specific context of our use cases (Sec. 3.3). We specifically focus on the trade-off between drone location privacy and utility, i.e., the extent to which drone-based LBSs can successfully use the disclosed location.
@@
Experiments Settings. To answer the RQs above, we designed a measurement system to obtain the execution time, RAM usage, communication overhead, energy consumption, location privacy and utility of our approach. As for the execution time, we measured the time (in milliseconds) necessary to generate a RID message secured through OLO-RID, from the acquisition of the current location from the GNSS (not included) to the assembling of the entire RID message (wireless communication not included). We specifically excluded location acquisition and wireless communication as they are orthogonal processes to our solution, which might be executed asynchronously. As for the RAM usage, we used the Linux tool /usr/bin/time -v, providing the maximum instantaneous RAM occupancy of a process. Through this tool, we can obtain an indication of the amount of RAM to be deployed on the involved devices to support our solution fully. As for the communication cost, using standard tools provided by default C libraries, we measured the size of the complete RID message after the application of our solution. We also measure the energy consumption (in milliJoule) the drone requires to run our solution. To this end, in line with similar contributions, we used the energy estimation tool powertop provided by the Linux OS. Powertop estimates the power (in Watts) of a running process based on hardware-dependent power consumption figures. To compute the corresponding energy consumption $E$, we multiply the obtained power estimate $P$ for the duration of the process of interest $\Delta T$ (the duration of the runtime phase), measured as described above, as $E=P\times\Delta T$. Such a value can indicate the energy required to run our solution and, thus, the battery to be installed on the drone to keep broadcasting (location) privacy-preserving RID messages for the whole mission. To evaluate the location privacy provided to the drone thanks to the application of our solution, in line with the relevant literature [11], we use the average distance, i.e., the average value of the (absolute) difference between the actual drone location and the disclosed (obfuscated) one. The higher the average distance, the more significant the uncertainty on the real drone’s location, thus, the higher the location privacy. Finally, as for the utility of the drone’s disclosed location, we ran simulations in Matlab using real drone traces (see below the reference dataset used in our analysis). We used metrics specific to the considered use case. For the use case #1, we considered the number of TPs (actual detected invasions of no-fly area), FPs (false detected invasions of no-fly area), TNs (correct identified non-invasions of the no-fly area) and FNs (missed detected invasions of the no-fly area) as seen by the invasion monitoring system. For the use case #2, we considered the average extra distance, i.e., the extra distance the drone has to travel to reach the suggested charging station compared to the distance to reach the optimal one. Finally, for the use case #3, we evaluate the average extra distance as the additional space the signal emitted from the suggested UAV has to travel to reach the user compared to the distance it would travel from the UAV closest to the requesting user.
[…]
@@
We ran all tests at least $5,000$ times, so as to average the results over multiple locations and trajectories. We also ran all tests using the highest privileges on the specific hardware, avoiding delays due to background concurrent processes. As a reference, we show in Fig. 7 the relationship between the actual drone path, the released obfuscated locations and the WiFi coverage (upper bound on the location privacy).
@@ 8.2 Experiment 1: Overhead Evaluation
As intuition would suggest, the average time required to run our solution increases with the size of the EC points (keys) used in the curve, i.e., with the provided security level. Considering the two curves NIST521 and BLS48556, both providing 256-bit security, we can see a slight discrepancy between the results obtained on the laptop and those obtained on the RPI. However, this difference is negligible ($3.5$), and it does not affect the rationale of our findings. It is essential to notice that all our results, both on the laptop and on the RPI, report execution times well below the threshold of $1$ s. The minimal execution times demonstrate the viability of our solution on relevant hardware.
@@
We notice that, on the particular hardware platform, the RAM consumption of our solution is independent of the chosen security level. On the laptop, our solution requires $3,600$ kB of RAM in the worst case, which is very lightweight and affordable even on low-end drones. We highlight that it is not possible to cross-compare the RAM consumption on the two different platforms because the underlying memory management system of such platforms is different.
Communication. To evaluate the communication overhead of OLO-RID, we can compare the regular RID format with the RID message format required by our solution, summarized in Table III. OLO-RID requires changing the content of three fields (current latitude, longitude, and altitude of the drone) and delivering three additional fields. The modified fields require the same amount of bytes as the regular fields, generating no additional overhead. As for the three added components, i.e., ephemeral key, ciphertext, and HMAC tag, the ciphertext and HMAC tag required by ECIES have a constant size of 16 and 32 bytes, respectively. The size of the ephemeral key depends on the adopted curve and, as per Tab. II), it varies between 64 and 140 bytes. Thus, the total communication overhead of OLO-RID for a single message ranges between $16+32+64=112$ Bytes and $16+32+140=188$ Bytes. The maximum RID message size with OLO-RID amounts to $227$ bytes, which is well below WiFi’s MTU, i.e., 2,304 bytes [21].
@@
In line with Fig. 8, increasing the number of security bits increases the energy consumption of our solution. The energy expenditure is primarily influenced by the time necessary to complete the protocol. We can notice the same difference between the curves NIST521 and BLS48556 among the two testbeds, as discussed above. The most energy-consuming curve on the laptop is NIST521, which requires $600$ mJ for protocol run. The most energy-consuming curve on the RPI is BLS548556, which requires an average of $1,100$ mJ per protocol run. We cannot cross-compare the energy expenditure across different hardware platforms, due to the different components and voltage needed to power them. We discuss the impact of the measured energy expenditure on an average drone in Sec. 9.
@@ 8.3 Experiment 2: Privacy Analysis
For answering RQ2, we analyze the impact of $\epsilon$ on the location privacy provided to the drone and the overhead incurred by OLO-RID. We recall from Sec. 5 that OLO-RID uses $\epsilon$ to tune the privacy level of the extended PIM obfuscation mechanism. Tab. IV reports the average distance of the disclosed location to the actual one for different values of the parameter $\epsilon$, with $\delta=0.01$.
@@
Changing $\epsilon$ affects the average distance significantly, from just a few meters ($2.54$ meters with $\epsilon=1$) to a few kilometres ($2,533.38$ meters with $\epsilon=0.001$). On the drone, lower values of $\epsilon$ provide higher average distance values and, thus, more privacy. However, we have to consider also the utility of the disclosed location to LBSs. We analyze the utility of the disclosed locations in Sec. 8.4 while, in the rest of this analysis, we focus on three reference values of $\epsilon$, i.e., $0.01,0.05,0.1$. Fig. 11 and Fig. 12 report the average time and the energy (with 95% confidence intervals) required to compute a RID message through our solution on the RPI with such values of $\epsilon$, respectively.
@@
From Fig. 11, we can see no impact on the execution time when we change $\epsilon$. We notice minor changes among the different $\epsilon$ values. In general, $\epsilon=0.1$ uses more energy than the others—in the worst case, $18$% more. When looking specifically at the higher security ECs $\epsilon=0.01$ is more energy efficient. Overall, the trend found in Fig. 8 still applies, i.e., the EC curve used as part of ECIES is the most relevant source of energy consumption in the protocol. We also notice oscillating results when investigating the protocol’s RAM consumption while varying $\epsilon$, as shown in Fig. 13. Here, we notice that choosing $\epsilon=0.1$ is always more memory-efficient than the other options, but with a small margin of $3$%, while $\epsilon=0.01$ and $\epsilon=0.05$ approximately use the same amount of RAM. As all of the values of $\epsilon$ report similar results, we believe the specific value of $\epsilon$ does not affect the RAM overhead of our solution significantly.
@@ 8.4 Experiment 3: Use Cases
Use Case 1: Invasion of No-Fly Zone. Consider the Use Case 1 described in Sec. 3.3 and particularly, the scenario depicted in Fig. 14, where the big red circle represents the CI location, the straight black path represents the NFZ boundaries, the dotted black path represents the WA bounds, and the blue path identifies the drone path.
Figure 14: Use Case # 1 analysis: sketch of the NFZ and WA surrounding a CI, with an actual drone flight path passing through.
Recall the definitions of TPs, FPs, TNs, and FNs given in Sec. 8.1. In general, we aim to have very high TPs and TNs, as well as very low FPs and FNs. However, as timely drone detection takes priority, we aim to keep the TPs high while reducing the FPs as much as possible. We consider two reference values for the average distance, i.e., 25 m and 100 m. We set the radius of the NFZ to 500 m, and we consider two values for the radius of the WA, i.e., 505 m and 600 m, to investigate to what extent the difference between the NFZ and the WA impacts our findings. Figs. 15 and 16 report the confusion matrices of the detection system considering the average distance of 25 m and the WA radius of 505 and 600 m, respectively.
@@
We notice that the size of the WA does not impact the TPs, while it does affect the FPs. In fact, increasing the radius of the WA, even though the percentage of FPs decreases, the actual number of FPs increases. Considering that each FP involves an interaction over the Internet between the CI observer and the USS (see Sec. 5.3), the CI operator can choose a small WA, characterized by a radius similar to the one of the NFZ, so as to reduce FPs and, accordingly, the required interactions with the USS. We can notice the trend described above also when comparing Figs. 17 and 18, considering the same analysis as above but with an average distance of $100$ m in the locations disclosed by the drone. Increasing the average distance to $100$ m increases the FNs significantly compared to using an average distance of $50$ meters, thus increasing the chances that the CI misses an invading drone. At the same time, with an average distance of $100$ m, the FPs also increase, as there are more chances that a drone which is not invading discloses a location inside the NFZ, thus falsely resulting as invading. To decrease the number of FPs, the WA should be as small as possible. While increasing the average distance increases FNs, there are more chances that the invasion is detected within a few seconds, resulting in a TP. Also, when the WA is small, increasing the average distance barely affects the number of FPs.
@@
Without OLO-RID, the system would suggest to the UAV the nearest charging station to the actual UAV location. With OLO-RID in place, as the UAV discloses an obfuscated location, the system suggests the charging station closest to the reported obfuscated location. The best case is when the suggested charging station coincides with the ones that the system would have suggested if the actual UAV location was disclosed, as the extra distance to be travelled by the UAV because of the application of location obfuscation is 00 m. When this is not the case, the extra distance the UAV has to travel should be as small as possible. In the following, we investigate the utility of OLO-RID in this use case by evaluating the extra distance travelled by the UAV to reach the suggested charging station with different protocol configurations. We consider a UAV disclosing locations with average distance values of 25 m, 50 m, and 100 m, and a variable number of deployed charging stations, i.e., 2, 8, and 32, so to investigate the impact of a variable density of such charging stations on the utility of our solution. We summarize the results of our analysis in Tab. V.
@@
Increasing the average distance of the locations disclosed by the drone (i.e., increasing drone location privacy) leads to an increase in the average extra distance, i.e., the charging station suggested by the system is not optimal. At the same time, increasing the number of charging stations deployed by the system also increases the average extra distance, as there is more chance that another charging station is closer to the disclosed location than to the actual location. Asymptotically, if we consider an area completely filled with charging stations, there would always be a charging station coincidental with the UAV’s disclosed location. Therefore, the average extra distance asymptotically increases to match the average distance value. To provide further insights, we show in Figs. 20 and 21 the probability distribution of the extra distance travelled by the UAV considering the average distance of $100$ m with 2 and 32 deployed charging stations, respectively.
@@
When the number of charging stations (density) increases, the probability of picking a suboptimal charging station also increases. We also notice that, excluding the values for the extra distance 00 m, the remaining values tend to converge to a Gaussian distribution, with mean values coincidental with the average distance. Such a finding confirms our intuition that, asymptotically, the extra distance coincides with the average distance.
Use Case 3: UAV as a service. Consider Use Case #3 described in Sec. 3.3, and particularly, the scenario depicted in Fig. 22, where the blue marker depicts a user requesting a service and the yellow lines indicate the flight path of various UAVs providing services.
@@
We notice that the results reported in Tab. VI for Use Case #3 align with the ones in Tab. V for Use Case #2. Increasing the average distance, i.e., increasing drones’ location privacy, leads to a higher extra distance, i.e., suboptimal performances of the location-based service. Similarly, increasing the number of UAVs increases the average extra distance, leading to suboptimal performances. We can apply the same reasoning as for Use Case #2 to reason about the asymptotic behavior of the extra distance in Use Case #3: assuming there is a UAV providing service on top of each possible location (i.e., an infinite number of UAVs), the average extra distance would match precisely the average distance. Finally, we notice a similarity between Use Case #2 and Use Case # 3 also when analyzing the probability distribution of the extra distance for specific configurations, as shown in Figs. 23 and 24 considering an average distance of $100$ m with 2 and 10 UAVs, respectively.
@@
When the UAVs providing services increase, the user’s probability of picking the suboptimal UAV also increases. Excluding the values for the extra distance 00 m, the remaining values tend to converge to a Gaussian distribution, with mean values coincidental with the value of the average distance on the drone. This finding aligns with our intuition that, asymptotically, the extra distance coincides with the average distance.

### v2

@@ 7 Proof-of-Concept Implementation
Hardware Details. We use two different devices as a UAV, i.e., a general-purpose laptop and an embedded device, to evaluate the performance of OLO-RID on high-end and medium-end UAVs. Our first PoC uses a general-purpose laptop equipped with one CPU running at 2.60GHz and 16GB of Random Access Memory (RAM). Such features align with the resources available onboard medium-high-end UAVs, e.g., the DJI Matrice 300 [djiMatrice]. The second PoC uses as the UAV a RPI Model 3B+, featuring a 1.4GHz CPU and 1GB of RAM [raspberry_pi]. The resources of the RPI align with the ones of medium-low-end UAVs such as the DJI Mini 2 [dji_mini_2]. Also, note that such a choice aligns with the methodology used by other scientific contributions in the UAV security domain [george2023_sacmat]. Both the laptop and the RPI integrate an IEEE 802.11b/g/n module, useful for transmitting and receiving RID-compliant WiFi messages over the air.
Software. We implemented OLO-RID using the C programming language to reduce runtime delays deriving from language interpretation as much as possible. We used the well-known library MIRACL CORE to implement cryptography operations required by the ECIES encryption scheme [miracl]. Note that we chose MIRACL Core due to its well-known excellent performance on low-powered IoT systems and the available support for several elliptic curves, as required for running ECIES. Such a choice also aligns with relevant literature on security solutions for constrained UAVs [wisse2023_iotj]. We report in Tab. II the Elliptic Curves we used within the ECIES scheme. To show the difference between curves of the same security levels, we have also considered two additional curves, namely BN254 and BLS48556. Using such curves could give us information about the difference in performance between two curves with the same level of security. Note that the curve BN254 was designed for 128-bit security [bn254] but, recently, its actual security decreased to 100-bit [bn254_issues].
@@
BN254 64 100 [bn254_issues]
secp256r1 (NIST256) 64 128 [nist_curves]
secp384r1 (NIST384) 96 192 [nist_curves]
secp521r1 (NIST521) 132 256 [nist_curves]
BLS48556 140 256 [bls48]
@@
We ran OLO-RID on the RPI as a native process. On the laptop, to isolate it from concurrent processes, we ran OLO-RID within a Linux-based virtual environment featuring 4 CPU cores and 8GB of RAM. The source code of our solution is available open-source at [code].
@@ 8.1 Experimental Methodology
  • RQ1: What is the runtime overhead of OLO-RID on regular UAVs regarding computation time, RAM usage, energy expenditure, and communication bandwidth with various security levels?
@@
  • RQ2: How can we improve the location privacy provided to the UAV at runtime, and how do such improvements affect the overhead of our solution?
@@
  • RQ3: What is the utility of the location disclosed through OLO-RID by the UAV across various configuration parameters while considering our reference use cases discussed in Sec. 3.3?
@@
UAVs are constrained devices, characterized by (one or more) constraints regarding available computational, memory, communication bandwidth, and energy resources. Such constraints make it challenging to implement and run a suitable technique for location obfuscation while still creating and broadcasting RID messages within the time limitations imposed by the RID regulation ($1$ second). RQ1 sets a baseline for the overhead of our solution when using different hardware and various ECs. RQ2 investigates the impact in terms of location privacy and overhead on the UAV derived by changing relevant location obfuscation parameters (i.e., $\epsilon$). Finally, RQ3 evaluates OLO-RID in the specific context of our use cases (Sec. 3.3). We specifically focus on the trade-off between UAV location privacy and utility, i.e., the extent to which UAV-based LBSs can successfully use the disclosed location.
@@
Experiments Settings. To answer the RQs above, we designed a measurement system to obtain the execution time, RAM usage, communication overhead, energy consumption, location privacy and utility of our approach. As for the execution time, we measured the time (in milliseconds) necessary to generate a RID message secured through OLO-RID, from the acquisition of the current location from the GNSS (not included) to the assembling of the entire RID message (wireless communication not included). We specifically excluded location acquisition and wireless communication as they are orthogonal processes to our solution, which might be executed asynchronously. As for the RAM usage, we used the Linux tool /usr/bin/time -v, providing the maximum instantaneous RAM occupancy of a process. Through this tool, we can obtain an indication of the amount of RAM to be deployed on the involved devices to support our solution fully. As for the communication cost, using standard tools provided by default C libraries, we measured the size of the complete RID message after the application of our solution. We also measure the energy consumption (in milliJoule) the UAV requires to run our solution. To this end, in line with similar contributions, we used the energy estimation tool powertop provided by the Linux OS. Powertop estimates the power (in Watts) of a running process based on hardware-dependent power consumption figures. To compute the corresponding energy consumption $E$, we multiply the obtained power estimate $P$ for the duration of the process of interest $\Delta T$ (the duration of the runtime phase), measured as described above, as $E=P\times\Delta T$. Such a value can indicate the energy required to run our solution and, thus, the battery to be installed on the UAV to keep broadcasting (location) privacy-preserving RID messages for the whole mission. To evaluate the location privacy provided to the UAV thanks to the application of our solution, in line with the relevant literature [brighente2022_ares], we use the average distance, i.e., the average value of the (absolute) difference between the actual UAV location and the disclosed (obfuscated) one. The higher the average distance, the more significant the uncertainty on the real UAV’s location, thus, the higher the location privacy. Finally, as for the utility of the UAV’s disclosed location, we ran simulations in Matlab using real UAV traces (see below the reference dataset used in our analysis). We used metrics specific to the considered use case. For the use case #1, we considered the number of TPs (actual detected invasions of no-fly area), FPs (false detected invasions of no-fly area), TNs (correct identified non-invasions of the no-fly area) and FNs (missed detected invasions of the no-fly area) as seen by the invasion monitoring system. For the use case #2, we considered the average extra distance, i.e., the extra distance the UAV has to travel to reach the suggested charging station compared to the distance to reach the optimal one. Finally, for the use case #3, we evaluate the average extra distance as the additional space the signal emitted from the suggested UAV has to travel to reach the user compared to the distance it would travel from the UAV closest to the user.
Reference Dataset. In this work, for flight data of the involved UAVs, we use the actual flight dataset provided by NATO’s Emerging Security Challenges Division [flight_data], in line with similar works [brighente2022_ares]. The data refer to multiple flights paths of various UAVs in an area of $1.5$ km $\times$ $2.6$ km $\times$ $40$ m. For all the tests, we use the flight path shown in Fig. 5, including over $5,000$ location disclosures in its flight path. Instead, for the analysis of the use cases, we use the flight path shown in Fig. 6, which includes over $8,000$ locations. We used a different flight path to analyze the use cases so as to work in a controlled environment, where the UAV flies mainly in a straight line. Such a trajectory eliminates anomalies due to weird flight paths and makes it easier to reason about the entire scenario.
@@
We ran all tests at least $5,000$ times, so as to average the results over multiple locations and trajectories. We also ran all tests using the highest privileges on the specific hardware, avoiding delays due to background concurrent processes. As a reference, we show in Fig. 7 the relationship between the actual UAV path, the released obfuscated locations and the WiFi coverage (upper bound on the location privacy).
@@ 8.2 Experiment 1: Overhead Evaluation
As intuition would suggest, the average time required to run our solution increases with the size of the EC points (keys) used in the curve, i.e., with the provided security level. Considering the two curves NIST521 and BLS48556, both providing 256-bit security, we can see a slight discrepancy between the results obtained on the laptop and those obtained on the RPI. However, this difference is negligible ($3.5$ %), and it does not affect the rationale of our findings. It is essential to notice that all our results, both on the laptop and on the RPI, report execution times well below the threshold of $1$ s. The minimal execution times demonstrate the viability of our solution on relevant hardware.
@@
We notice that, on the particular hardware platform, the RAM consumption of our solution is independent of the chosen security level. On the laptop, our solution requires $3,600$ kB of RAM in the worst case, which is very lightweight and affordable even on low-end UAVs. We highlight that it is not possible to cross-compare the RAM consumption on the two different platforms because the underlying memory management system of such platforms is different.
Communication. To evaluate the communication overhead of OLO-RID, we can compare the regular RID format with the RID message format required by our solution, summarized in Table III. OLO-RID requires changing the content of three fields (current latitude, longitude, and altitude of the UAV) and delivering three additional fields. The modified fields require the same amount of bytes as the regular fields, generating no additional overhead. As for the three added components, i.e., ephemeral key, ciphertext, and HMAC tag, the ciphertext and HMAC tag required by ECIES have a constant size of 16 and 32 bytes, respectively. The size of the ephemeral key depends on the adopted curve and, as per Tab. II), it varies between 64 and 140 bytes. Thus, the total communication overhead of OLO-RID for a single message ranges between $16+32+64=112$ Bytes and $16+32+140=188$ Bytes. The maximum RID message size with OLO-RID amounts to $227$ bytes, which is well below WiFi’s MTU, i.e., 2,304 bytes [ieee2012].
@@
In line with Fig. 8, increasing the number of security bits increases the energy consumption of our solution. The energy expenditure is primarily influenced by the time necessary to complete the protocol. We can notice the same difference between the curves NIST521 and BLS48556 among the two testbeds, as discussed above. The most energy-consuming curve on the laptop is NIST521, which requires $600$ mJ for protocol run. The most energy-consuming curve on the RPI is BLS548556, which requires an average of $1,100$ mJ per protocol run. We cannot cross-compare the energy expenditure across different hardware platforms, due to the different components and voltage needed to power them. We discuss the impact of the measured energy expenditure on an average UAV in Sec. 9.
@@ 8.3 Experiment 2: Privacy Analysis
For answering RQ2, we analyze the impact of $\epsilon$ (privacy budget) on the location privacy provided to the UAV and the overhead incurred by OLO-RID. We recall from Sec. 5 that OLO-RID uses $\epsilon$ to tune the privacy level of the extended PIM obfuscation mechanism. Tab. IV reports the average distance of the disclosed location to the actual one for different values of the parameter $\epsilon$, with $\delta=0.01$.
@@
Changing $\epsilon$ affects the average distance significantly, from just a few meters ($2.54$ meters with $\epsilon=1$) to a few kilometres ($2,533.38$ meters with $\epsilon=0.001$). On the UAV, lower values of $\epsilon$ provide higher average distance values and, thus, more privacy. However, we have to consider also the utility of the disclosed location to LBSs. We analyze the utility of the disclosed locations in Sec. 8.4 while, in the rest of this analysis, we focus on three reference values of $\epsilon$, i.e., $0.01,0.05,0.1$. Fig. 11 and Fig. 12 report the average time and the energy (with 95% confidence intervals) required to compute a RID message through our solution on the RPI with such values of $\epsilon$, respectively.
@@
From Fig. 11, we can see no impact on the execution time when we change $\epsilon$. We notice minor changes among the different $\epsilon$ values. In general, $\epsilon=0.1$ uses more energy than the others—in the worst case, 1818% more. When looking specifically at the higher security ECs $\epsilon=0.01$ is more energy efficient. Overall, the trend found in Fig. 8 still applies, i.e., the EC curve used as part of ECIES is the most relevant source of energy consumption in the protocol. We also notice oscillating results when investigating the protocol’s RAM consumption while varying $\epsilon$, as shown in Fig. 13. Here, we notice that choosing $\epsilon=0.1$ is always more memory-efficient than the other options, but with a small margin of 33%, while $\epsilon=0.01$ and $\epsilon=0.05$ approximately use the same amount of RAM. As all of the values of $\epsilon$ report similar results, we believe the specific value of $\epsilon$ does not affect the RAM overhead of our solution significantly.
@@ 8.4 Experiment 3: Use Cases
Use Case 1: Invasion of No-Fly Zone. Consider the Use Case 1 described in Sec. 3.3 and particularly, the scenario depicted in Fig. 14, where the big red circle represents the CI location, the straight black path represents the NFZ boundaries, the dotted black path represents the WA bounds, and the blue path identifies the UAV path.
Figure 14: Use Case # 1 analysis: sketch of the NFZ and WA surrounding a CI, with an actual UAV flight path passing through.
Recall the definitions of TPs, FPs, TNs, and FNs given in Sec. 8.1. In general, we aim to have very high TPs and TNs, as well as very low FPs and FNs. However, as timely UAV detection takes priority, we aim to keep the TPs high while reducing the FPs as much as possible. We consider two reference values for the average distance, i.e., 25 m and 100 m. We set the radius of the NFZ to 500 m, and we consider two values for the radius of the WA, i.e., 505 m and 600 m, to investigate to what extent the difference between the NFZ and the WA impacts our findings. Figs. 15 and 16 report the confusion matrices of the detection system considering the average distance of 25 m and the WA radius of 505 and 600 m, respectively.
@@
We notice that the size of the WA does not impact the TPs, while it does affect the FPs. In fact, increasing the radius of the WA, even though the percentage of FPs decreases, the actual number of FPs increases. Considering that each FP involves an interaction over the Internet between the CI observer and the USS (see Sec. 5.3), the CI operator can choose a small WA, characterized by a radius similar to the one of the NFZ, so as to reduce FPs and, accordingly, the required interactions with the USS. We can notice the trend described above also when comparing Figs. 17 and 18, considering the same analysis as above but with an average distance of $100$ m in the locations disclosed by the UAV. Increasing the average distance to $100$ m increases the FNs significantly compared to using an average distance of $50$ meters, thus increasing the chances that the CI misses an invading UAV. At the same time, with an average distance of $100$ m, the FPs also increase, as there are more chances that a UAV which is not invading discloses a location inside the NFZ, thus falsely resulting as invading. To decrease the number of FPs, the WA should be as small as possible. While increasing the average distance increases FNs, there are more chances that the invasion is detected within a few seconds, resulting in a TP. Also, when the WA is small, increasing the average distance barely affects the number of FPs.
@@
Without OLO-RID, the system would suggest to the UAV the nearest charging station to the actual UAV location. With OLO-RID in place, as the UAV discloses an obfuscated location, the system suggests the charging station closest to the reported obfuscated location. The best case is when the suggested charging station coincides with the ones that the system would have suggested if the actual UAV location was disclosed, as the extra distance to be travelled by the UAV because of the application of location obfuscation is 0 m. When this is not the case, the extra distance the UAV has to travel should be as small as possible. In the following, we investigate the utility of OLO-RID in this use case by evaluating the extra distance travelled by the UAV to reach the suggested charging station with different protocol configurations. We consider a UAV disclosing locations with average distance values of 25 m, 50 m, and 100 m, and a variable number of deployed charging stations, i.e., 2, 8, and 32, so to investigate the impact of a variable density of such charging stations on the utility of our solution. We summarize the results of our analysis in Tab. V.
@@
Increasing the average distance of the locations disclosed by the UAV (i.e., increasing UAV location privacy) leads to an increase in the average extra distance, i.e., the charging station suggested by the system is not optimal. At the same time, increasing the number of charging stations deployed by the system also increases the average extra distance, as there is more chance that another charging station is closer to the disclosed location than to the actual location. Asymptotically, if we consider an area completely filled with charging stations, there would always be a charging station coincidental with the UAV’s disclosed location. Therefore, the average extra distance asymptotically increases to match the average distance value. To provide further insights, we show in Figs. 20 and 21 the probability distribution of the extra distance travelled by the UAV considering the average distance of $100$ m with 2 and 32 deployed charging stations, respectively.
@@
When the number of charging stations (density) increases, the probability of picking a suboptimal charging station also increases. We also notice that, excluding the values for the extra distance 0 m, the remaining values tend to converge to a Gaussian distribution, with mean values coincidental with the average distance. Such a finding confirms our intuition that, asymptotically, the extra distance coincides with the average distance.
Use Case 3: UAV-as-a-service. Consider Use Case #3 described in Sec. 3.3, and particularly, the scenario depicted in Fig. 22, where the blue marker depicts a user requesting a service and the yellow lines indicate the flight path of various UAVs providing services.
@@
We notice that the results reported in Tab. VI for Use Case #3 align with the ones in Tab. V for Use Case #2. Increasing the average distance, i.e., increasing UAVs’ location privacy, leads to a higher extra distance, i.e., suboptimal performances of the location-based service. Similarly, increasing the number of UAVs increases the average extra distance, leading to suboptimal performances. We can apply the same reasoning as for Use Case #2 to reason about the asymptotic behavior of the extra distance in Use Case #3: assuming there is a UAV providing service on top of each possible location (i.e., an infinite number of UAVs), the average extra distance would match precisely the average distance. We provide more insights into the service degradation caused by privacy preservation by reporting the additional attenuation incurred by a sample signal in an exemplary deployment. We consider a transmitter emitting signals at the frequency of 2.5 GHz, located at random coordinates of latitude 42.3001 and longitude -71.3504. We also considered a random receiver, operating on the same frequency, located at latitude 42.3467 and longitude -71.0972. For such an example deployment, we consider the propagation model provided by Matlab for heavy rain, and we compute the path loss at the receiver site using such a propagation model. Then, to show the impact of our solution, we computed the path loss experienced by the receiver when considering the average extra distance in the worst case of Tab. VI, i.e., 100 m of average distance in the locations disclosed by the UAV and a number of 10 service UAVs. Without applying our solution, the path loss experienced by the receiver is 127.320793 db. Such a path loss becomes 127.320796 db at an average extra distance of 1.19 6m (average distance of 25 m in the locations disclosed by the UAVs), 127.320801 db at an average extra distance of 4.46 m (average distance of 50 m in the locations disclosed by the UAVs), and finally 127.320822 db at an average extra distance of 16.636 m (average distance of 100 m in the locations disclosed by the UAVs). Therefore, even in most privacy-preserving scenario (100 meters of average distance in the locations disclosed by the UAVs), the users experience a degradation of the received signal of only 0.000029 db, which is clearly very limited, thus confirming the value of our solution. Finally, we notice a similarity between Use Case #2 and Use Case # 3 also when analyzing the probability distribution of the extra distance for specific configurations, as shown in Figs. 23 and 24 considering an average distance of $100$ m with 2 and 10 UAVs, respectively.
@@
When the UAVs providing services increase, the user’s probability of picking the suboptimal UAV also increases. Excluding the values for the extra distance 0 m, the remaining values tend to converge to a Gaussian distribution, with mean values coincidental with the value of the average distance on the UAV. This finding aligns with our intuition that, asymptotically, the extra distance coincides with the average distance.